# Jet Tagging with More-Interaction Particle Transformer

### v1

@@ IV Conclusion
  • On the Top Tagging Dataset: The MIParT model achieved accuracy and AUC metrics nearly identical to those of LorentzNet, and its ${\rm Rej}_{50\%}$ and ${\rm Rej}_{30\%}$ metrics are comparable within the error range to LorentzNet. The MIParT model significantly outperformed ParT in the top tagging benchmark, achieving approximately 25% better background rejection at a 30% signal efficiency. Among the models evaluated, MIParT, along with LorentzNet, ranks at the top tier, showcasing some of the most robust performances.
  • On the Quark-gluon Dataset: The MIParT model significantly outperforms LorentzNet across all metrics, including accuracy, AUC, ${\rm Rej}_{50\%}$, and ${\rm Rej}_{30\%}$, as well as several other models. MIParT achieved the best performance across all evaluation metrics, improving background rejection power by approximately 3% compared to ParT.
Overall, MIParT outperformed ParT on both the top tagging and quark-gluon tagging tasks while also exhibiting lower computational complexity and fewer parameters. Previously, it was generally assumed that transformer-based models required large-scale dataset pretraining to achieve optimal results. Our MIParT model demonstrates that with higher-dimensional particle interaction embeddings, top-tier performance can be achieved without pretraining on large datasets, even surpassing ParT.
Furthermore, while pretraining ParT on the larger multi-class JETCLASS dataset and subsequently fine-tuning it on the top tagging dataset can enhance performance, we plan to explore this approach for MIParT in future work. Moreover, MIParT is not limited to jet tagging but can also be applied to event identification, which could be immensely helpful in the search for new physics signals.
@@ Acknowledgements.
The authors would like to thank Huilin Qu for helpful discussions. This work was supported by the National Natural Science Foundation of China (NNSFC) under grant Nos. 12275066 and 11605123.

### v2

@@ IV Conclusion
  • On the Top Tagging Dataset: The MIParT model achieved accuracy and AUC metrics nearly identical to those of LorentzNet, and its ${\rm Rej}_{50\%}$ and ${\rm Rej}_{30\%}$ metrics are comparable within the error range to LorentzNet. And a series of Lorentz-equivariant methods demonstrated performance similar to that of LorentzNet. The MIParT model significantly outperformed ParT in the top tagging benchmark, achieving approximately 25% better background rejection at a 30% signal efficiency. Among the models evaluated, MIParT, along with LorentzNet and several other Lorentz-equivariant based models, ranks at the top tier, showcasing some of the most robust performances. For the fine-tuned MIParT-L model that is pre-trained on the 100M JetClass dataset, a 39% enhancement in background rejection performance was achieved, comparable to that of the fine-tuned ParT.
  • On the Quark-gluon Dataset: The MIParT model significantly outperforms LorentzNet across all metrics, including accuracy, AUC, ${\rm Rej}_{50\%}$, and ${\rm Rej}_{30\%}$, as well as several other models. MIParT achieved the best performance across all evaluation metrics, improving background rejection power by approximately 3% compared to ParT. For the fine-tuned MIParT-L model, background rejection performance improved by 6%, surpassing that of the fine-tuned ParT. Specifically, the background rejection of fine-tuned MIParT-L improved by an additional 2% compared to the fine-tuned ParT.
Overall, MIParT outperformed ParT on both the top tagging and quark-gluon tagging tasks while also exhibiting lower computational complexity and fewer parameters. Previously, it was generally assumed that transformer-based models required large-scale dataset pre-training to achieve optimal results. Our MIParT model demonstrates that with higher-dimensional particle interaction embeddings, top-tier performance can be achieved without pre-training on large datasets, even surpassing ParT.
Furthermore, as pre-training ParT on the larger multi-class JetClass dataset and subsequently fine-tuning it on the top tagging dataset can enhance performance, we have applied this approach to MIParT-L in this work. We find that MIParT-L can further capitalize on the knowledge from large datasets, showing superior capabilities after fine-tuning. Specifically, it performs better on the quark-gluon dataset than the fine-tuned ParT. Finding more efficient ways to fine-tune a base Transformer model will be especially helpful for future experiments when generic and foundation models are deployed, and downstream application tasks are varied. Moreover, MIParT is not limited to jet tagging but can also be applied to event identification, which could be immensely helpful in the search for new physics signals.
@@ Acknowledgements.
The work of K. Wang and J. Zhu is supported by the National Natural Science Foundation of China (NNSFC) under grant Nos. 12275066 and 11605123.